Is Z[sqrt(-5)] factorial ring?
1
Expert's answer
2012-08-09T08:12:21-0400
Element 6=2*3=(1+i*sqrt(5))(1-i*sqrt(5)).
We have found element with non-unique decomposition, so it is not factorial.